8x-6y=17
y=4/3x+2

Answer by HyperBrain(694)

Simply substitute y=4/3x+2 to 8x-6y=17:
8x-6(4/3x+2)=17
8x-(8x+12)=17
-12=17
---
oops! We obtained a false statement.
---
Therefore, there is no solution.